OpenStudy (anonymous):

How do I solve: -2= 6/7p Thanks :) P.S will give medals <3

OpenStudy (shane_b):

Do you have an idea of where to start?

OpenStudy (shane_b):

Also, is it \[-2=\frac{6}{7}p\]OR\[-2=\frac{6}{7p}\]

OpenStudy (lgbasallote):

^couldnt resist the bribe of medals? lol jk

OpenStudy (anonymous):

I have no idea where to start??

OpenStudy (shane_b):

Ok...which of the versions I posted is correct?

OpenStudy (anonymous):

version 1

OpenStudy (anonymous):

Please help!!!

OpenStudy (anonymous):

ok so this is your equation: \[-2=\frac{ 6 }{ 7 }p\] what u have to do is multiply by the reciprocal on both sides of the equation so that p will become 1 \[\frac{ 7 }{ 6 }\times -2=\frac{ 7 }{ 6 }\times \frac{ 6 }{ 7 }p\] so \[\frac{ 7 }{ 6}\times \frac{ 6 }{ 7 }\] is cancelled and your left with \[\frac{ 7 }{ 6 }\times \frac{ 2 }{ 1 }=p\] now u simplify by dividing 6 by 2 and 2 by 2then you'll have \[\frac{ 7 }{ 3}\times \frac{ 1 }{ 1}=p\] then u multiply and you end up with \[p =\frac{ 7 }{ 3 }\]

OpenStudy (shane_b):

I gave up when there was no reply for a while :)

OpenStudy (anonymous):

lol it just takes time

OpenStudy (shane_b):

I was answering, not asking...I get impatient at nearly 6am!